Operation Summary for May 25, 2011 - May 27, 2011 |
|
Wednesday May 25 The day shift began with Operations monitoring the Tevatron (TEV) store 8765 (116.09E30) and the Recycler stash (281.3E10), with the Antiproton Source (Pbar) stacking (15.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, TEV personnel began crystal collimator studies at 10 AM and the studies ran throughout the shift. Linac personnel conducted some destructive beam studies in the 750 KeV line. Jasmin studiers made several accesses into the upper Pbar vault. Operators searched and secured MI-65 and then NuMI personnel began conducting horn scans. RF personnel worked on two MI RF stations (MIRF6 and 7). During the evening shift, Operations terminated store 8765 at 4:36 PM. TEV personnel completed their studies at 4:40 PM. Operations put the TEV into HEP shot setup and established store 8766 at 6:01 PM with an average initial luminosity (AIL) of 383.17E30. NuMI personnel completed their horn scans at 10:30 PM; operators prepared MI-65 for access. The Pbar AP50 chilled water water skid tripped off; operators reset it. Operators tuned the Linac beam. Thursday May 26 The midnight shift began with Operations monitoring store 8766 (150.94E30) and the stash (192.8E10), with Pbar stacking (24.2E10), and with MiniBooNE taking beam. During the midnight shift, NuMI personnel reported a PLC problem with their compressed air system at 6:38 AM, which will require an outside contractor to repair. Operators searched and secured MTA. During the day shift, Operations terminated store 8766 at 8:30 AM, put the TEV into shot setup, and established store 8767 at 9:38 AM with an AIL of 386.5E30. Operators search and secured the Meson MTest enclosures and then began sending beam to the returning FTBF experiment T-978 (CALICE) at 3:28 PM. During the evening shift, Operations disabled beam to T-978 at 6:04 PM. At 10:46 PM, the Recycler Pelletron suffered a “LOSTI” trip (a fluctuation of the dome current); operators reset it. At 11:06 PM, a Pbar Debuncher Lambertson (D:ELAM) power supply appeared to have tripped off, but it had only toggled into local mode. Operators resolved the problem. D0 reported a problem with their primary dry engine. Cryo system personnel began preparing the backup engine for operation. Friday May 27 The midnight shift began with Operations monitoring store 8767 (82.86E30) and the stash (379.2E10), with Pbar stacking (24.2E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8767 at 12:49 AM, put the TEV into shot setup, and established store 8768 at 1:59 AM with an AIL of 405.42E30. D0 reported at 3:55 AM that cryo system personnel had the backup dry engine on and were beginning repairs on the primary engine. Operations established beam to T-978 at 5:15 AM. The Plans for Friday and the Weekend The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to conduct a study. |
|
More Information Comments and Suggestions |
| last modified 05/27/2011 email Fermilab |
| Security, Privacy, Legal | |